# A preorder on the set of links with applications to symmetric unions

### v2

@@ Abstract.
For a link $L$ in the $3$-sphere, the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ is defined as a quotient of the link group of $L$. When there exists an epimorphism $G^{\mathrm{orb}}(L)\to G^{\mathrm{orb}}(L^{\prime})$, we denote this by $L\succeq L^{\prime}$ and explore the relationships between the two links. Specifically, we prove that if $L\succeq L^{\prime}$ and $L$ is a Montesinos link with $r$ rational tangles $(r\geq 3)$, then $L^{\prime}$ is either a Montesinos link with at most $r+1$ rational tangles or a certain connected sum. We further show that if $L$ is a small link, then there are only finitely many links $L^{\prime}$ satisfying $L\succeq L^{\prime}$. In contrast, if $L$ has determinant zero, then $L\succeq L^{\prime}$ for every $2$-bridge link $L^{\prime}$. Our main applications concern symmetric unions of knots. In particular, we provide a criterion showing that a given knot does not admit a symmetric union presentation.
@@ 1. Introduction
$1\to\pi_{1}(\Sigma_{2}(L))\to G^{\mathrm{orb}}(L)\to\mathbb{Z}/2\mathbb{Z}\to 1.$
In particular, if $\Sigma_{2}(L)$ is aspherical, the only torsion elements in $G^{\mathrm{orb}}(L)$ come from the meridians and have order $2$.
@@ 1.1. A preorder defined via orbifolds
$L_{1}$ $\pi$-dominates $L_{2}$, written $L_{1}\succeq L_{2}$, if there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L_{1})\twoheadrightarrow G^{\mathrm{orb}}(L_{2})$.
It is worth noting that a symmetric union construction gives plenty of examples of $L_{1}\succeq L_{2}$ (see Section 7). By [9] and [12, Theorem 3.1], the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ determines the pair $(S^{3},L)$ up to homeomorphism when $L$ is a prime link with at least three bridges. The $3$-manifold group $\pi_{1}(\Sigma_{2}(L))$ is residually finite by [31] and of index $2$ in $G^{\mathrm{orb}}(L)$. It follows that the group $G^{\mathrm{orb}}(L)$ is a (finitely generated) residually finite group, and hence hopfian. Then the relation $\succeq$ is a partial order on the set of prime links with at least three bridges, up to mirror image. The anti-symmetry is not obvious, but it follows from the facts that $\pi$-orbifold groups are hopfian and that prime links with at least three bridges are determined by their $\pi$-orbifold groups.
Here are some straightforward properties from the definition.
Proposition 1.2.
Let $L$ be a link in $S^{3}$.
  1. (1)
  2. (2)
    Any link $\pi$-dominates the unknot.
  3. (3)
    A connected sum $L_{1}\sharp L_{2}$ $\pi$-dominates each summand $L_{1}$ and $L_{2}$.
  4. (4)
    A link $L$ $\pi$-dominates each sublink $L^{\prime}\subset L$, in particular, each component.
  5. (5)
@@ 1.4. Symmetric unions
Finally, we use the $\pi$-orbifold group to study a classical construction in knot theory introduced by Kinoshita and Terasaka in [39] and generalized by Lamm [42], which always produces ribbon knots. This construction is a generalization of the connected sum of a knot with its mirror image and called a symmetric union. It gives a $\pi$-domination between the symmetric union and the partial knot (see Section 7). Hence, our results (Theorems 1.5, 1.8 and 1.11) provide strong constraints for a knot to be the partial knot of a symmetric union (see Corollaries 7.3, 7.4, and 7.5). These tools can be applied to the open problem of whether every ribbon knot admits a symmetric union presentation.
@@ Acknowledgments
This study was supported in part by JSPS KAKENHI Grant Numbers JP19K03505, JP23K12974, and JP24H00686. The first and second authors were supported by Soka University International Collaborative Research Grant.
@@ 2.1. Montesinos links
First, we recall fundamental properties of a Montesinos link. Let $L=L(\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ be a Montesinos link with $r\geq 3$ rational tangles of slopes $\beta_{i}/\alpha_{i}\in\mathbb{Q}$ with $\alpha_{i}>1$ and $\beta_{i}\neq 0$ coprime to $\alpha_{i}$. It has been shown by Montesinos [48] that the $2$-fold branched cover $\Sigma_{2}(L)$ of $L$ is the closed orientable Seifert fibered $3$-manifold $V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ with base a $2$-dimensional orientable orbifold with underlying space $S^{2}$ and $r$ singular points with branching indices $\alpha_{i}\geq 2$, corresponding to the $r$ exceptional fibers of types $(\alpha_{i},\beta_{i})$. See also [17, Chapter 12.D]. Its rational Euler number $e_{0}=\sum_{i=1}^{r}\frac{\beta_{i}}{\alpha_{i}}\in\mathbb{Q}$ verifies that
@@ 2.2. Orbifolds
Assume $L\succeq L^{\prime}$, that is, there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$. There is a meridian $\mu\in G^{\mathrm{orb}}(L)$ whose image $\varphi(\mu)$ does not belong to the normal subgroup $\pi_{1}(\Sigma_{2}(L^{\prime}))$ of index $2$, otherwise $\varphi(G^{\mathrm{orb}}(L))=G^{\mathrm{orb}}(L^{\prime})$ would be in $\pi_{1}(\Sigma_{2}(L^{\prime}))$. It follows that $\varphi$ sends the subgroup $\pi_{1}(\Sigma_{2}(L))$ of index $2$ onto the subgroup $\pi_{1}(\Sigma_{2}(L^{\prime}))$ of index $2$, and therefore $\varphi$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Now, the proof of Theorem 1.5 uses the orbifold theorem, see [9, Theorem 1].
@@
The center $Z=\langle h\rangle$ of $\pi_{1}(\Sigma_{2}(L))$ is a infinite cyclic normal subgroup in $G^{\mathrm{orb}}(L)$ generated by an element $h$ such that $\bar{\mu}h\bar{\mu}^{-1}=h^{-1}$, where $\bar{\mu}$ is the image of a meridian $\mu$ of $G(L)$ in the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$. This follows from the fact that, for a Montesinos link, the covering involution $\tau$ on $\Sigma_{2}(L)$ is fiber-preserving and reverses the orientation of the fibers (see [48, Section 2] and [12, Theorem 1.3]). The image $\varphi(Z)$ of the center $Z$ of $\pi_{1}(\Sigma_{2}(L))$ is in the center of $\pi_{1}(\Sigma_{2}(L^{\prime}))$.
@@
(II) Assume that $L^{\prime}$ is not prime and has at most three bridges. First, we consider the case where $L^{\prime}$ is split: $L^{\prime}=L^{\prime}_{1}\sqcup L^{\prime}_{2}$. Then $L^{\prime}$ cannot be the $2$-component trivial link $U\sqcup U$ which is a $2$-bridge link. It can always be expressed as a non-trivial connected sum $L_{1}\sharp(U\sqcup U)\sharp L_{2}$. It follows from Lemma 3.1(1) and (2) that $\Sigma_{2}(L^{\prime})=\Sigma_{2}(L^{\prime}_{1})\sharp(S^{1}\times S^{2})\sharp\Sigma_{2}(L^{\prime}_{1})$, with $\Sigma_{2}(L^{\prime}_{1})$ or $\Sigma_{2}(L^{\prime}_{2})$ not being $S^{3}$. Therefore, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\mathbb{Z}\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{2}))$ is a non-trivial free product, and hence centerless. The epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$ must kill the infinite cyclic center $Z$ of $\pi_{1}(\Sigma_{2}(L))$. Thus $\tilde{\varphi}$ factors through $\Gamma=\pi_{1}(\Sigma_{2}(L))/Z$ and induces an epimorphism onto each factor of $\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\mathbb{Z}\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{2}))$. As mentioned at the beginning of (3), $\Gamma$ is generated by torsion elements, and thus each factor must contain some torsion element. Therefore, $L$ cannot be split.
@@
We now assume that $L^{\prime}$ is not split. Since $L^{\prime}$ is not prime by Theorem 2.2, $L^{\prime}$ decomposes into a connected sum $L^{\prime}=L^{\prime}_{1}\sharp\cdots\sharp L^{\prime}_{n}$ of finitely many prime links which are unique up to permutations ($n\geq 2$). Then, by Lemma 3.1(1)–(3), $\Sigma_{2}(L^{\prime})\cong\Sigma_{2}(L^{\prime}_{1})\sharp\cdots\sharp\Sigma_{2}(L^{\prime}_{n})$ is a non-trivial connected sum of closed, orientable, irreducible $3$-manifolds. In particular, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\cdots\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{n}))$ is a non-trivial free product and hence centerless. The argument given above for the split case shows that each factor $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ must contain some torsion element ($1\leq i\leq n$). Since $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ is the fundamental group of a closed irreducible orientable $3$-manifold for $1\leq i\leq n$, it must be of finite order by [32, Corollary 9.9]. By the orbifold theorem, each $\Sigma_{2}(L^{\prime}_{i})$ is a Seifert $3$-manifold with finite fundamental group, and each link $L^{\prime}_{i}$ is either a $2$-bridge link or an elliptic Montesinos link with three rational tangles. Let $n_{1}$ be the number of 2-bridge factors of $L^{\prime}$ and $n_{2}$ the number of elliptic Montesinos factors of $L^{\prime}$. Then, the number $n=n_{1}+n_{2}$ of prime factors of $L^{\prime}$ verifies $n_{1}+2n_{2}\leq r-1$ since the fundamental group of the $2$-fold branched cover of a $2$-bridge link has rank $1$, that of an elliptic Montesinos link has rank $2$, and that of a Montesinos link with $r\geq 3$ rational tangles has rank at most r1r-1.
@@
Let $N$ be a closed orientable irreducible $3$-manifold whose JSJ pieces are all hyperbolic. Then $N$ is aspherical and thus $\pi_{1}(N)$ is torsion-free. Moreover, by [34, Theorem VI.1.6(i)], the centralizer in $\pi_{1}(N)$ of any non-trivial element is isomorphic to $\mathbb{Z}$ or $\mathbb{Z}\oplus\mathbb{Z}$ (see also [26]), hence abelian. Then, Theorem 1.8 follows from Theorem 4.7 and Corollary 4.10, where the proof of Theorem 4.7 is based on Lemmas 4.1 and 4.3 below.
@@
Since $M_{0}$ is a rational homology solid torus, one can choose a simple closed curve $\alpha\subset\partial M_{0}$ such that the Dehn filling $M_{0}(\alpha)$ is a rational homology sphere (see [62, Lemma 3.2] for example). The natural epimorphism
@@
Using Corollary 4.10, we prove Theorem 1.8 concerning an arborescent link $L$ with $\det L\neq 0$.
@@
Let $L$ be an arborescent link with $\det L\neq 0$ and suppose $L\succeq L^{\prime}$. If $L^{\prime}$ is not prime, then it is a connected sum $L_{1}\sharp L_{2}$ or a split union $L_{1}\sqcup L_{2}$, and thus we have $L\succeq L_{1}$ and $L\succeq L_{2}$. It follows from Lemma 3.1(2) that $\Sigma_{2}(L_{1}\sharp L_{2})=\Sigma_{2}(L_{1})\sharp\Sigma_{2}(L_{2})$ and $\Sigma_{2}(L_{1}\sqcup L_{2})=\Sigma_{2}(L_{1})\sharp(S^{1}\times S^{2})\sharp\Sigma_{2}(L_{2})$. Therefore, it suffices to prove Theorem 1.8 when $L^{\prime}$ is prime, which is equivalent to $\Sigma_{2}(L^{\prime})$ being irreducible by Lemma 3.1(3). The epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $\Sigma_{2}(L)$ is a graph manifold which is a rational homology sphere and $\Sigma_{2}(L^{\prime})$ is irreducible, Corollary 4.10 implies that a JSJ piece of $\Sigma_{2}(L^{\prime})$ must be Seifert fibered. ∎

### v3

@@ Abstract.
For a link $L$ in the $3$-sphere, the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ is defined as a quotient of the link group $G(L)$ of $L$. When there exists an epimorphism $G^{\mathrm{orb}}(L)\to G^{\mathrm{orb}}(L^{\prime})$ fitting into a certain commutative diagram, we define a relation $L\succeq L^{\prime}$ and explore the relationships between the two links. Specifically, we prove that if $L\succeq L^{\prime}$ and $L$ is a Montesinos link with $r$ rational tangles $(r\geq 3)$, then $L^{\prime}$ is either a Montesinos link with at most $r+1$ rational tangles or a certain connected sum. We further show that if $L$ is a small link, then there are only finitely many links $L^{\prime}$ satisfying $L\succeq L^{\prime}$. In contrast, if $L$ has determinant zero, then $L\succeq L^{\prime}$ for every $2$-bridge link $L^{\prime}$. Our main applications concern symmetric unions of knots. In particular, we provide a criterion showing that a given knot does not admit a symmetric union presentation.
@@ 1. Introduction
$1\to\pi_{1}(\Sigma_{2}(L))\to G^{\mathrm{orb}}(L)\xrightarrow{\nu}\mathbb{Z}/2\mathbb{Z}\to 1,$
where the epimorphism $\nu$ sends the image of each meridian to the generator of $\mathbb{Z}/2\mathbb{Z}$. If $\Sigma_{2}(L)$ is aspherical, $\pi_{1}(\Sigma_{2}(L))$ is torsion-free and the only torsion elements in $G^{\mathrm{orb}}(L)$ come from the meridians, and hence they are of order $2$.
@@ 1.1. A preorder defined via orbifolds
$L$ $\pi$-dominates $L^{\prime}$, written $L\succeq L^{\prime}$, if there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ fitting into the commutative diagram
${G^{\mathrm{orb}}(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$$\scriptstyle{\varphi}$${\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${G^{\mathrm{orb}}(L^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{\prime}}$${\mathbb{Z}/2\mathbb{Z}.}$
See Section 2.3 for basic properties of the $\pi$-domination. It is worth noting that a symmetric union construction gives plenty of examples of pairs $(K,K^{\prime})$ of knots such that $K\succeq K^{\prime}$ (see Section 7). By [9] and [12, Theorem 3.1], the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$ determines the pair $(S^{3},L)$ up to homeomorphism when $L$ is a prime link with at least three bridges. The $3$-manifold group $\pi_{1}(\Sigma_{2}(L))$ is residually finite by [31] and of index $2$ in $G^{\mathrm{orb}}(L)$. It follows that the group $G^{\mathrm{orb}}(L)$ is a (finitely generated) residually finite group, and hence hopfian. Then the relation $\succeq$ is a partial order on the set of prime links with at least three bridges, up to mirror image. The reflexivity and transitivity follow immediately from the definition. The anti-symmetry is not obvious, but it follows from the facts that $\pi$-orbifold groups are hopfian and that prime links with at least three bridges are determined by their $\pi$-orbifold groups.
@@ 1.4. Symmetric unions
Finally, we use the $\pi$-orbifold group to study a classical construction in knot theory introduced by Kinoshita and Terasaka in [39] and generalized by Lamm [42], which always produces ribbon knots. This construction is a generalization of the connected sum of a knot with its mirror image and called a symmetric union. It gives a $\pi$-domination between the symmetric union and the partial knot (see Section 7). Hence, our results (Theorems 1.4, 1.7, and 1.10) provide strong constraints for a knot to be the partial knot of a symmetric union (see Corollaries 7.3, 7.4, and 7.5). These tools can be applied to the open problem of whether every ribbon knot admits a symmetric union presentation.
@@ Acknowledgments
This study was supported in part by JSPS KAKENHI Grant Numbers JP19K03505, JP23K12974, and JP24H00686. The first and second authors were supported by Soka University International Collaborative Research Grant. The authors are grateful to the anonymous referee for crucial comments improving the manuscript.
@@ 2.1. Montesinos links
First, we recall fundamental properties of a Montesinos link. Let $L=L(\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ be a Montesinos link with $r\geq 3$ rational tangles of slopes $\beta_{i}/\alpha_{i}\in\mathbb{Q}$ with $\alpha_{i}>1$ and $\beta_{i}\neq 0$ coprime to $\alpha_{i}$. It has been shown by Montesinos [48] that the $2$-fold branched cover $\Sigma_{2}(L)$ of $L$ is the closed orientable Seifert fibered $3$-manifold $V(0;e_{0};\frac{\beta_{1}}{\alpha_{1}},\dots,\frac{\beta_{r}}{\alpha_{r}})$ with base a $2$-dimensional orientable orbifold with underlying space $S^{2}$ and $r$ singular points with branching indices $\alpha_{i}\geq 2$, corresponding to the $r$ exceptional fibers of types $(\alpha_{i},\beta_{i})$. See also [17, Chapter 12.D]. Its rational Euler number $e_{0}=\sum_{i=1}^{r}\frac{\beta_{i}}{\alpha_{i}}\in\mathbb{Q}$ satisfies
@@ 2.2. Orbifolds
2.3. $\pi$-domination
We recall the definition of $\pi$-domination between links in $S^{3}$. Let $L$ and $L^{\prime}$ be links in $S^{3}$, $L$ $\pi$-dominates $L^{\prime}$, written $L\succeq L^{\prime}$, if there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ such that the following diagram is commutative:
${G^{\mathrm{orb}}(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu}$$\scriptstyle{\varphi}$${\mathbb{Z}/2\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{id}}$${G^{\mathrm{orb}}(L^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu^{\prime}}$${\mathbb{Z}/2\mathbb{Z}.}$
Here, we show some basic properties of $\pi$-domination.
Proposition 2.9.
Let $L$ be a link in $S^{3}$.
  1. (1)
  2. (2)
    Any link $\pi$-dominates the unknot.
  3. (3)
    A connected sum $L_{1}\sharp L_{2}$ $\pi$-dominates each summand $L_{1}$ and $L_{2}$.
  4. (4)
Proof.
For a link $L\subset S^{3}$, the abelianization $G^{\mathrm{orb}}(L)^{\mathrm{ab}}$ of the $\pi$-orbifold group is isomorphic to $(\mathbb{Z}/2\mathbb{Z})^{\#L}$. Then, (1) follows from the fact that $\varphi$ induces an epimorphism between the abelianizations of the $\pi$-orbifold groups. Moreover, (2) is immediate from the epimorphism $G^{\mathrm{orb}}(L)\xrightarrow{\nu}\mathbb{Z}/2\mathbb{Z}\cong G^{\mathrm{orb}}(U)$, where $U$ denotes the unknot.
(3) For a connected sum, we have $G^{\mathrm{orb}}(L_{1}\sharp L_{2})=G^{\mathrm{orb}}(L_{1})\ast_{\mathbb{Z}/2\mathbb{Z}}G^{\mathrm{orb}}(L_{2})$. Using the epimorphism $G^{\mathrm{orb}}(L_{2})\xrightarrow{\nu}\mathbb{Z}/2\mathbb{Z}$, one gets an epimorphism $\varphi\colon G^{\mathrm{orb}}(L_{1}\sharp L_{2})\to G^{\mathrm{orb}}(L_{1})\ast_{\mathbb{Z}/2\mathbb{Z}}\mathbb{Z}/2\mathbb{Z}=G^{\mathrm{orb}}(L_{1})$, which fits into the diagram of Definition 1.1.
(4) There is an epimorphism
$G^{\mathrm{orb}}(L_{1}\sqcup L_{2})=G^{\mathrm{orb}}(L_{1})\ast G^{\mathrm{orb}}(L_{2})\to G^{\mathrm{orb}}(L_{1})\ast_{\mathbb{Z}/2\mathbb{Z}}G^{\mathrm{orb}}(L_{2})=G^{\mathrm{orb}}(L_{1}\sharp L_{2})$
which amalgamates the two cyclic subgroups of order $2$ generated by the images of the meridians involved in the connected sum. This epimorphism fits into the diagram of Definition 1.1. ∎
The next lemma shows that the relation $\succeq$ is strongly related to the domination relation on the set of closed orientable $3$-manifolds defined via epimorphisms between their fundamental groups.
Lemma 2.10.
Let $L$ and $L^{\prime}$ be links in $S^{3}$. If $L\succeq L^{\prime}$ via an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$, then $\varphi$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$.
Proof.
We have the commutative diagram
[…]
where the two rows are exact. It follows from a diagram chasing that $\varphi$ induces a homomorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\to\pi_{1}(\Sigma_{2}(L^{\prime}))$ fitting into the above diagram. Furthermore, $\tilde{\varphi}$ is surjective by a diagram chasing. ∎
The next lemma states that, for a given epimorphism between $\pi$-orbifold groups, if the domain arises from a knot, then the commutativity of the diagram in Definition 1.1 always holds. This remains true for links $L$ and $L^{\prime}$ with the same number of components when $\Sigma_{2}(L^{\prime})$ is aspherical.
Lemma 2.11.
  1. (1)
    If $K$ is a knot and there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(K)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$, then $L^{\prime}$ is a knot and $K\succeq L^{\prime}$.
  2. (2)
    Suppose that $L=\#L^{\prime}$ and that $\Sigma_{2}(L^{\prime})$ is aspherical. If there is an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$, then $L\succeq L^{\prime}$.
Proof.
(1) Since $G^{\mathrm{orb}}(K)^{\mathrm{ab}}\cong\mathbb{Z}/2\mathbb{Z}$, $L^{\prime}$ is a knot by the proof of Proposition 2.9(1). Let $\bar{\mu}\in G^{\mathrm{orb}}(K)$ be the image of any meridian $\mu\in G(K)$. Since $\varphi(\bar{\mu})$ normally generates $G^{\mathrm{orb}}(L^{\prime})$, $\varphi(\bar{\mu})$ does not belong to the normal subgroup $\pi_{1}(\Sigma_{2}(L^{\prime}))$ of index $2$ in $G^{\mathrm{orb}}(L^{\prime})$, and thus $\nu^{\prime}\circ\varphi(\bar{\mu})\neq 0$.
(2) The induced epimorphism $\bar{\varphi}\colon G^{\mathrm{orb}}(L)^{\mathrm{ab}}\to G^{\mathrm{orb}}(L^{\prime})^{\mathrm{ab}}$ is an isomorphism since the abelian group $(\mathbb{Z}/2\mathbb{Z})^{\#L}$ is hopfian. Then, for $\bar{\mu}\in G^{\mathrm{orb}}(L)$, $\varphi(\bar{\mu})$ is of order $2$. Hence, it does not belong to $\pi_{1}(\Sigma_{2}(L^{\prime}))$ which is torsion-free by the hypothesis. Therefore, $\varphi(\bar{\mu})$ maps to the generator of the quotient $G^{\mathrm{orb}}(L^{\prime})/\pi_{1}(\Sigma_{2}(L^{\prime}))\cong\mathbb{Z}/2\mathbb{Z}$. ∎
@@ 2.3. $\pi$-domination
Assume $L\succeq L^{\prime}$ via an epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$. Then, by Lemma 2.10, $\varphi$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Now, the proof of Theorem 1.4 uses the orbifold theorem, see [9, Theorem 1].
@@
The center $Z=\langle h\rangle$ of $\pi_{1}(\Sigma_{2}(L))$ is an infinite cyclic normal subgroup in $G^{\mathrm{orb}}(L)$ generated by an element $h$ such that $\bar{\mu}h\bar{\mu}^{-1}=h^{-1}$, where $\bar{\mu}$ is the image of a meridian $\mu$ of $G(L)$ in the $\pi$-orbifold group $G^{\mathrm{orb}}(L)$. This follows from the fact that, for a Montesinos link, the covering involution $\tau$ on $\Sigma_{2}(L)$ is fiber-preserving and reverses the orientation of the fibers (see [48, Section 2] and [12, Theorem 1.3]). The image $\varphi(Z)$ of the center $Z$ of $\pi_{1}(\Sigma_{2}(L))$ is in the center of $\pi_{1}(\Sigma_{2}(L^{\prime}))$.
@@
(II) Assume that $L^{\prime}$ is not prime and has at most three bridges. First, we consider the case where $L^{\prime}$ is split: $L^{\prime}=L^{\prime}_{1}\sqcup L^{\prime}_{2}$. Then $L^{\prime}$ cannot be the $2$-component trivial link $U\sqcup U$ which is a $2$-bridge link. It can always be expressed as a non-trivial connected sum $L^{\prime}_{1}\sharp(U\sqcup U)\sharp L^{\prime}_{2}$. It follows from Lemma 3.1(1) and (2) that $\Sigma_{2}(L^{\prime})=\Sigma_{2}(L^{\prime}_{1})\sharp(S^{1}\times S^{2})\sharp\Sigma_{2}(L^{\prime}_{2})$, with $\Sigma_{2}(L^{\prime}_{1})$ or $\Sigma_{2}(L^{\prime}_{2})$ not being $S^{3}$. Therefore, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\mathbb{Z}\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{2}))$ is a non-trivial free product, and hence centerless. The epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$ must kill the infinite cyclic center $Z$ of $\pi_{1}(\Sigma_{2}(L))$. Thus $\tilde{\varphi}$ factors through $\Gamma=\pi_{1}(\Sigma_{2}(L))/Z$ and induces an epimorphism onto each factor of $\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\mathbb{Z}\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{2}))$. As mentioned at the beginning of (3), $\Gamma$ is generated by torsion elements, and thus each factor must contain some torsion element. Therefore, $L$ cannot be split.
@@
We now assume that $L^{\prime}$ is not split. Since $L^{\prime}$ is not prime by Theorem 2.2, $L^{\prime}$ decomposes into a connected sum $L^{\prime}=L^{\prime}_{1}\sharp\cdots\sharp L^{\prime}_{n}$ of finitely many prime links which are unique up to permutations ($n\geq 2$). Then, by Lemma 3.1(1)–(3), $\Sigma_{2}(L^{\prime})\cong\Sigma_{2}(L^{\prime}_{1})\sharp\cdots\sharp\Sigma_{2}(L^{\prime}_{n})$ is a non-trivial connected sum of closed, orientable, irreducible $3$-manifolds. In particular, $\pi_{1}(\Sigma_{2}(L^{\prime}))=\pi_{1}(\Sigma_{2}(L^{\prime}_{1}))\ast\cdots\ast\pi_{1}(\Sigma_{2}(L^{\prime}_{n}))$ is a non-trivial free product and hence centerless. The argument given above for the split case shows that each factor $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ must contain some torsion element ($1\leq i\leq n$). Since $\pi_{1}(\Sigma_{2}(L^{\prime}_{i}))$ is the fundamental group of a closed irreducible orientable $3$-manifold for $1\leq i\leq n$, it must be of finite order by [32, Corollary 9.9]. By the orbifold theorem, each $\Sigma_{2}(L^{\prime}_{i})$ is a Seifert $3$-manifold with finite fundamental group, and each link $L^{\prime}_{i}$ is either a $2$-bridge link or an elliptic Montesinos link with three rational tangles. Let $n_{1}$ be the number of 2-bridge factors of $L^{\prime}$ and $n_{2}$ the number of elliptic Montesinos factors of $L^{\prime}$. Then, the number $n=n_{1}+n_{2}$ of prime factors of $L^{\prime}$ satisfies $n_{1}+2n_{2}\leq r-1$ since the fundamental group of the $2$-fold branched cover of a $2$-bridge link has rank $1$, that of an elliptic Montesinos link has rank $2$, and that of a Montesinos link with $r\geq 3$ rational tangles has rank at most r1r-1.
@@
Let $N$ be a closed orientable irreducible $3$-manifold whose JSJ pieces are all hyperbolic. Then $N$ is aspherical and thus $\pi_{1}(N)$ is torsion-free. Moreover, by [34, Theorem VI.1.6(i)], the centralizer in $\pi_{1}(N)$ of any non-trivial element is isomorphic to $\mathbb{Z}$ or $\mathbb{Z}\oplus\mathbb{Z}$ (see also [26]), hence abelian. Then, Theorem 1.7 follows from Theorem 4.7 and Corollary 4.10, together with Lemma 2.10. The proof of Theorem 4.7 is based on Lemmas 4.1 and 4.3 below.
@@
Since $M_{0}$ is a rational homology solid torus, one can choose a simple closed curve $\alpha\subset\partial M_{0}$ such that the Dehn filling $M_{0}(\alpha)$ is a rational homology sphere (see [63, Lemma 3.2] for example). The natural epimorphism
@@
Using Corollary 4.10 and Lemma 2.10, we prove Theorem 1.7 concerning an arborescent link $L$ with $\det L\neq 0$.
@@
Let $L$ be an arborescent link with $\det L\neq 0$ and suppose $L\succeq L^{\prime}$. If $L^{\prime}$ is not prime, then it is a connected sum $L_{1}\sharp L_{2}$ or a split union $L_{1}\sqcup L_{2}$, and thus we have $L\succeq L_{1}$ and $L\succeq L_{2}$. It follows from Lemma 3.1(2) that $\Sigma_{2}(L_{1}\sharp L_{2})=\Sigma_{2}(L_{1})\sharp\Sigma_{2}(L_{2})$ and $\Sigma_{2}(L_{1}\sqcup L_{2})=\Sigma_{2}(L_{1})\sharp(S^{1}\times S^{2})\sharp\Sigma_{2}(L_{2})$. Therefore, it suffices to prove Theorem 1.7 when $L^{\prime}$ is prime, which is equivalent to $\Sigma_{2}(L^{\prime})$ being irreducible by Lemma 3.1(3). By Lemma 2.10, the epimorphism $\varphi\colon G^{\mathrm{orb}}(L)\twoheadrightarrow G^{\mathrm{orb}}(L^{\prime})$ induces an epimorphism $\tilde{\varphi}\colon\pi_{1}(\Sigma_{2}(L))\twoheadrightarrow\pi_{1}(\Sigma_{2}(L^{\prime}))$. Since $\Sigma_{2}(L)$ is a graph manifold which is a rational homology sphere and $\Sigma_{2}(L^{\prime})$ is irreducible, Corollary 4.10 implies that a JSJ piece of $\Sigma_{2}(L^{\prime})$ must be Seifert fibered. ∎